# Exploring Radar Data Representations in Autonomous Driving: A Comprehensive Review

### v2

@@ III-D3 Discussion
However, as grid maps are constructed from radar point clouds, point cloud sparsity impacts detection, segmentation, and localization accuracy. This may reduce the accuracy of fine-grained object details or small objects in the grid map. Moreover, closely spaced objects may be difficult to distinguish in radar grid maps.
@@ III-E1 Datasets
The Open Radar Datasets [67] collects different types of outdoor moving targets, such as pedestrian walking and cycling, captured by a stationary radar system. The primary goal of this dataset is to utilize classification techniques for distinguishing between different motion activities. Apart from outdoor environments, some micro-Doppler signature-based datasets (e.g., CI4R [66], Dop-NET [65], MCD-Gesture [68]) for motion recognition (e.g., health monitoring, gesture recognition) are from indoor environments. These indoor datasets can assist in occupant behavior recognition to some extent, but still differ from behavior classification in autonomous driving due to inconsistent acquisition environments.
@@ III-E3 Discussion
However, the transformation to the frequency domain via the Time-Frequency transform suffers from overlapping target components, resulting in suboptimal feature extraction for downstream tasks [75]. Similarly to ADC signal representation, the micro-Doppler signature is insufficient in delineating an object’s spatial coordinates for detection and segmentation tasks, so it is commonly employed for classification tasks. Moreover, the micro-Doppler signature representation may have limited resolution in range estimation compared to other representations like the radar tensor. This limitation could impact the ability to accurately determine the exact distance of an object, potentially affecting collision avoidance algorithms.
@@ IV-A2 Consideration of Computational Efficiency
If real-time processing and low computational complexity are critical, computationally efficient representations such as radar tensors and point clouds are preferable. The radar tensor represents radar echoes in a 2D, 3D or 4D matrix, while the point cloud representation is a set of individual 3D points representing detected objects. As far as computational performance is concerned, some tensor-based methods and point cloud-based methods have already achieved real-time object detection and segmentation [4, 170, 171, 20].
Conversely, ADC signals provide raw radar data with high temporal resolution, but are typically large and computationally expensive to process in real-time, requiring additional algorithms for signal processing and feature extraction [15, 81]. Micro-Doppler signatures may require additional processing, but their feasibility for real-time depends on specific algorithms and computational resources [172, 173]. Grid maps are structured grid-based formats, simplifying subsequent processing and potentially enabling computational efficiency. However, they may not meet real-time requirements if frequent updates are needed for new scenarios [174, 154].
@@ IV-B2 Dealing with Inherent Limitations
While combining multiple frames enhances accuracy, it can also introduce system delays [113, 181, 182, 133, 183, 58, 184]. Column or pillar expansion methods (e.g., [181, 182, 185]) complement the lack of measurement in the vertical direction, but only alleviate ambiguity to a limited extent. Exploring radar output distribution (e.g., Gaussian distribution) based on the fundamental principles of radar presents a valuable research opportunity [186].
Moreover, existing methods (e.g., PointNet [130], PointNet++ [131]) designed for processing dense LiDAR point clouds are limited for handling sparse radar point clouds [187]. Therefore, developing specialized architectures to address the challenges of sparse radar detection presents a promising research direction. For example, Liu et al[20] leveraged cosine similarity loss and normalized inner product loss as part of the training process for sparse radar detection points, enhancing the performance of point-wise center shift vector-guided clustering.
@@
For noise removal in radar tensors, researchers may draw inspiration from principles and techniques used in image denoising and restoration [192, 193, 194, 195, 196, 197, 198]. By considering radar tensors as images, various techniques, including convolutional neural networks, restoration filters, and wavelet-based methods, can be applied to reduce noise and enhance the quality of these tensors [199, 200]. For mitigating noise in point clouds, removing radar detections that fall outside boundaries defined by ground truth information from cameras or LiDARs is an approach [181, 201]. Constructing features based on physical properties can provide the necessary guidance for models to differentiate targets from clutter. This involves examining point density, spatial arrangement, and other statistical properties to help distinguish signal from noise. However, a unique challenge presented by radar point clouds is their inherent sparsity and inaccuracy. Developing techniques to account for this uncertainty could allow for more robust noise removal and aid in distinguishing objects from noise.
@@ IV-B3 Improving Data Processing Algorithms
Radar signals often exhibit complex patterns and non-linear dependencies. Traditional FFT-based methods may only capture part of the relevant information required for parameter estimation [204, 4]. One potential research opportunity is employing neural networks to extract radar parameters instead of traditional FFT operations. This approach is valuable to reduce computational requirements, while simplifying the data flow within embedded implementations [205, 4, 206]. An example can be found in RODNet [4], where FFT operations are selectively applied in the sample and antenna dimensions, while the chirp dimension is preserved to form the range-azimuth-chirp tensor. Then, a neural network is employed to process the chirp dimension for extracting relevant Doppler features. This end-to-end learning approach allows the neural network to learn relevant features and patterns from the radar data, potentially improving the accuracy and efficiency of parameter extraction.
@@ IV-B4 Employing 4D Radar Sensors
IV-C What about Data Fusion for Radar Representations?
While radar sensors are good at detecting and ranging objects, they provide limited information about object characteristics like color, texture, and fine-grained visual details. Integrating radars with complementary sensors, such as cameras or LiDARs, overcomes this limitation and provides comprehensive perception solutions [1, 221]. However, radar data fusion also faces some challenges, particularly in heterogeneous data association and in challenging condition processing. An overview of these challenges and research directions is summarized in Table V.
@@ IV-C1 Associating Heterogeneous Data
A significant challenge is the ambiguity in associating radar data with other modalities, as they are heterogeneous. For radar-camera fusion, existing approaches involve projecting radar data onto the image plane and subsequently establishing their correspondence via a calibration matrix [222, 108, 181, 223]. For radar-LiDAR fusion, existing methods exploit geometric information to associate radar and LiDAR point clouds by comparing the position, velocity, and shape of the detected objects to determine if they correspond to the same physical object [224]. However, projection-based or geometric-based associations result in ineffective alignment with object centers. As previously stated, radar data exhibits sparsity, inaccuracy, and noise, thereby causing inadequate associations in both object-level and data-level fusion scenarios.
@@ IV-C2 Handling Challenging Conditions
Fusion in challenging conditions may require adaptive fusion strategies that dynamically adjust fusion parameters based on environmental conditions. Developing data weighting techniques that assign appropriate weights to sensor measurements based on reliability can enhance fusion performance under challenging conditions [238, 239]. For example, radar sensors may be more reliable than cameras and LiDARs under adverse weather conditions. Therefore, the fusion algorithm should adaptively emphasize radar data and de-emphasize other sensors.
@@ Acknowledgment
This research was funded by Suzhou Municipal Key Laboratory for Intelligent Virtual Engineering (SZS2022004), Research Development Fund of XJTLU (RDF-19-02-23), XJTLU AI University Research Centre, Jiangsu Province Engineering Research Centre of Data Science and Cognitive Computation at XJTLU, SIP AI innovation platform (YZCXPT2022103) and Suzhou Science and Technology Project (SYG202122). This work received financial support from Jiangsu Industrial Technology Research Institute (JITRI) and Wuxi National Hi-Tech District (WND).

### v3

@@ III-D3 Discussion
However, as grid maps are constructed from radar point clouds, point cloud sparsity impacts detection, segmentation, and localization accuracy. This may reduce the accuracy of fine-grained object details or small objects in the grid map. Moreover, closely spaced objects are difficult to distinguish in radar grid maps.
@@ III-E1 Datasets
The Open Radar Datasets [70] collects different types of outdoor moving targets, such as pedestrian walking and cycling, captured by a stationary radar system. The primary goal of this dataset is to utilize classification techniques for distinguishing between different motion activities. Apart from outdoor environments, some micro-Doppler signature-based datasets (e.g., Dop-NET [69], MCD-Gesture [71]) for motion recognition (e.g., health monitoring, gesture recognition) are from indoor environments. These indoor datasets can assist in occupant behavior recognition to some extent, but still differ from behavior classification in autonomous driving due to inconsistent acquisition environments.
@@ III-E3 Discussion
However, the transformation to the frequency domain via the Time-Frequency transform suffers from overlapping target components, resulting in suboptimal feature extraction for downstream tasks [78]. Similarly to ADC signal representation, the micro-Doppler signature is insufficient in delineating an object’s spatial coordinates for detection and segmentation tasks, so it is commonly employed for classification tasks. Moreover, the micro-Doppler signature representation has limited resolution in range estimation compared to other representations like the radar tensor. This limitation could impact the ability to accurately determine the exact distance of an object, potentially affecting collision avoidance algorithms.
@@ IV-A1 Consideration of Application Scenarios
Autonomous driving vehicles are categorized into six levels of autonomy by the SAE International framework, ranging from Level 0 (no automation) to Level 5 (full automation) [171]. At lower levels of autonomy (Levels 1 and 2), traditional 3D radar sensors are often sufficient for tasks such as adaptive cruise control and lane-keeping assistance, where limited spatial resolution is acceptable. However, as vehicles progress to higher levels of autonomy (Levels 3 and 4), the demand for more precise and robust sensor data increases, making 4D radars a critical component due to their ability to provide denser point clouds, higher resolution, and accurate object tracking. At Level 5, fully autonomous systems rely on a fusion of sensors, including 4D radars, LiDAR, and cameras, to achieve robust perception in complex and dynamic environments.
@@ IV-A2 Consideration of Computational Efficiency
If real-time processing and low computational complexity are critical, computationally efficient representations such as radar tensors and point clouds are preferable. The radar tensor represents radar echoes in a 2D or 3D matrix, while the point cloud representation is a set of individual 3D points representing detected objects. As far as computational performance is concerned, some tensor-based methods and point cloud-based methods have already achieved real-time object detection and segmentation [4, 175, 176, 23].
Conversely, ADC signals or 4D radar tensor provide radar data with high temporal resolution, but are typically large and computationally expensive to process in real-time, requiring additional algorithms for signal processing and feature extraction [18, 84]. As mentioned in RADIal [42], the 4D radar tensor at each timestamp amounts to 450 MB, with a processing requirement of 45 GFLOPS for a single elevation, which increases to 495 GFLOPS when all 11 elevation are covered. On the other hand, the point cloud approach requires only 8 GFLOPS to compute the 3D coordinates for a sparse cloud of about 1000 points. In comparison, the processing of a 256-pixel image consumes only 0.4 GFLOPS of computational resources.
Micro-Doppler signatures require additional processing, but their feasibility for real-time depends on specific algorithms and computational resources [177, 178]. Grid maps are structured grid-based formats, simplifying subsequent processing and potentially enabling computational efficiency. However, they may not meet real-time requirements if frequent updates are needed for new scenarios [179, 158].
@@ IV-B2 Dealing with Inherent Limitations
The frequency of radar systems plays a critical role in determining data representation parameters, particularly resolution and the ability to discriminate between surfaces and objects with varying textures or roughness. Modern automotive radars, which operate at higher frequencies such as 77 GHz or 79 GHz, benefit from increased bandwidth, leading to improved range resolution. This allows for better differentiation between closely spaced objects and enhances the detection of subtle variations in distance. In addition to spatial resolution, higher frequencies enhance the radar’s ability to detect variations in surface textures and material properties, aiding in object classification and recognition. However, these advantages come with trade-offs, such as increased signal attenuation due to atmospheric absorption and reduced penetration through certain materials, which must be carefully considered in system design.
While combining multiple frames enhances accuracy, it can also introduce system delays [116, 186, 187, 137, 188, 61, 189]. Column or pillar expansion methods (e.g., [186, 187, 190]) complement the lack of measurement in the vertical direction, but only alleviate ambiguity to a limited extent. Exploring radar output distribution (e.g., Gaussian distribution) based on the fundamental principles of radar presents a valuable research opportunity [191]. Moreover, existing methods (e.g., PointNet [134], PointNet++ [135]) designed for processing dense LiDAR point clouds are limited for handling sparse radar point clouds [192]. Therefore, developing specialized architectures to address the challenges of sparse radar detection presents a promising research direction. For example, Liu et al[23] leveraged cosine similarity loss and normalized inner product loss as part of the training process for sparse radar detection points, enhancing the performance of point-wise center shift vector-guided clustering.
While conventional radar systems with a limited number of antennas may exhibit constrained angular resolution, this limitation can be effectively addressed through Synthetic Aperture Radar (SAR) techniques. By exploiting the motion of the vehicle, SAR can synthesize a large virtual aperture through advanced signal processing techniques, thereby enhancing angular resolution. This approach has been extensively studied in recent works [193, 194, 195, 196, 197, 198, 199], demonstrating its feasibility and effectiveness in automotive radar applications. For instance, [193] and [194] explored cooperative imaging in multi-sensor networks, particularly addressing phase synchronization and data processing to achieve high-resolution imaging. [195, 196, 197] analyzed the potential and challenges of automotive SAR imaging, proposing motion estimation and compensation techniques and navigation-aided imaging methods to mitigate the impact of vehicle motion on imaging quality. [198] and [199] analyzed the impact of urban environments on SAR imaging, proposing multi-beam SAR imaging techniques to address multi-path effects and occlusion issues in complex urban settings.
Given SAR’s ability to achieve spatial resolution comparable to LiDAR, the representation and storage of SAR data require careful consideration. High-resolution SAR data can be represented as point clouds, range-Doppler maps, or grid maps, with point clouds being particularly suitable for applications like autonomous driving due to their compatibility with LiDAR-based systems. However, the increased data volume necessitates efficient storage solutions, such as compression algorithms and hierarchical data structures. The choice of output format depends on the specific application, with point clouds being preferred for environment mapping and range-Doppler maps for tasks like target detection.
@@
For noise removal in radar tensors, researchers draw inspiration from principles and techniques used in image denoising and restoration [204, 205, 206, 207, 208, 209, 210]. By considering radar tensors as images, various techniques, including convolutional neural networks, restoration filters, and wavelet-based methods, can be applied to reduce noise and enhance the quality of these tensors [211, 212]. For mitigating noise in point clouds, removing radar detections that fall outside boundaries defined by ground truth information from cameras or LiDARs is an approach [186, 213]. Constructing features based on physical properties can provide the necessary guidance for models to differentiate targets from clutter. This involves examining point density, spatial arrangement, and other statistical properties to help distinguish signal from noise. However, a unique challenge presented by radar point clouds is their inherent sparsity and inaccuracy. Developing techniques to account for this uncertainty could allow for more robust noise removal and aid in distinguishing objects from noise.
@@ IV-B3 Improving Data Processing Algorithms
Radar signals often exhibit complex patterns and non-linear dependencies. Traditional FFT-based methods only capture part of the relevant information required for parameter estimation [216, 4]. One potential research opportunity is employing neural networks to extract radar parameters instead of traditional FFT operations. This approach is valuable to reduce computational requirements, while simplifying the data flow within embedded implementations [217, 4, 218]. An example can be found in RODNet [4], where FFT operations are selectively applied in the sample and antenna dimensions, while the chirp dimension is preserved to form the range-azimuth-chirp tensor. Then, a neural network is employed to process the chirp dimension for extracting relevant Doppler features. This end-to-end learning approach allows the neural network to learn relevant features and patterns from the radar data, potentially improving the accuracy and efficiency of parameter extraction.
@@ IV-B4 Employing 4D Radar Sensors
IV-C How Can Data Fusion Be Applied to Different Radar Representations?
While radar sensors are good at detecting and ranging objects, they provide limited information about object characteristics like color, texture, and fine-grained visual details. Integrating radars with complementary sensors, such as cameras, LiDARs and ultrasonic sensors, overcomes this limitation and provides comprehensive perception solutions [1, 233, 234, 235, 236]. Researchers can leverage the complementary strengths of each modality to achieve a balance between performance, cost, and robustness. Sensor fusion techniques further enhance system reliability by combining range and velocity data from radar with high-resolution spatial data from other sensors.
When fusing radar from multiple sensors, the choice of representation depends on the specific application requirements, the type of radar systems involved, and the desired level of information granularity. Common approaches to multi-radar fusion can be categorized into three levels: data-level fusion, feature-level fusion, and object-level fusion [1]. Data-level fusion involves combining the unprocessed radar signals from multiple radars. This approach preserves the maximum amount of information, as it avoids any loss of detail during intermediate processing steps. Feature-level fusion extracts meaningful features from the raw data of each radar system, such as target range, velocity, Doppler shift, and angular position. These features are then combined to form a unified representation of the detected targets. This approach strikes a balance between computational efficiency and information preservation, making it a popular choice for many practical applications. Object-level fusion involves processing the raw data from each radar system independently to make local decisions (e.g., object detection, classification, or tracking) and then combining these decisions to produce a final result. Common methods for object-level fusion include majority voting, weighted averaging, and Bayesian inference. However, radar data fusion also faces some challenges, particularly in heterogeneous data association and in challenging condition processing. An overview of these challenges and research directions is summarized in Table V.
@@ IV-C1 Associating Heterogeneous Data
A significant challenge is the ambiguity in associating radar data with other modalities, as they are heterogeneous. For radar-camera fusion, existing approaches involve projecting radar data onto the image plane and subsequently establishing their correspondence via a calibration matrix [237, 111, 186, 238, 16]. For radar-LiDAR fusion, existing methods exploit geometric information to associate radar and LiDAR point clouds by comparing the position, velocity, and shape of the detected objects to determine if they correspond to the same physical object [239]. However, projection-based or geometric-based associations result in ineffective alignment with object centers. As previously stated, radar data exhibits sparsity, inaccuracy, and noise, thereby causing inadequate associations in both object-level and data-level fusion scenarios.
@@ IV-C2 Handling Challenging Conditions
Fusion in challenging conditions requires adaptive fusion strategies that dynamically adjust fusion parameters based on environmental conditions. Developing data weighting techniques that assign appropriate weights to sensor measurements based on reliability can enhance fusion performance under challenging conditions [253, 254]. For example, radar sensors may be more reliable than cameras and LiDARs under adverse weather conditions. Therefore, the fusion algorithm should adaptively emphasize radar data and de-emphasize other sensors.
@@ Acknowledgment
This research was funded by Suzhou Municipal Key Laboratory for Intelligent Virtual Engineering (SZS2022004), Research Development Fund of XJTLU (RDF-19-02-23), XJTLU AI University Research Centre, Jiangsu Province Engineering Research Centre of Data Science and Cognitive Computation at XJTLU and SIP AI innovation platform (YZCXPT2022103). This work received financial support from Jiangsu Industrial Technology Research Institute (JITRI) and Wuxi National Hi-Tech District (WND).